OpenStudy (anonymous):

Find the exact value by using a half-angle identity sin7pi/8

OpenStudy (anonymous):

Any help is appreciated thanks

OpenStudy (anonymous):

(sin 7pi/8)^2=(1-cos7pi/4)/2

OpenStudy (anonymous):

the answers are -1/2sqrt1-sqrt2 1/2sqrt1-sqrt2 1/2sqrt2-sqrt2 -1/2sqrt2-sqrt2

OpenStudy (anonymous):

cos7pi/4=cospi/4=sqrt2 /2

OpenStudy (anonymous):

..so combine the two.and solve it.the ans is near to0.38